OpenStudy (anonymous):

A town has a population of 32,000 in the year 2002; 38,400 in the year 2003; 46,080 in the year 2004; and 55,296 in the year 2005. If this pattern continues, what will the population be in the year 2015?

OpenStudy (anonymous):

285,315?

OpenStudy (anonymous):

im figuring out the problem right now.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

OOpz i already figured out :P = 55,296(1.2)^10 = 55,296(6.1917364224) = 342,378

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Thank you